INFRARED SPECTRA OF N_{2}-BROADENED ^{13}CH_{4} AT TITAN ATMOSPHERIC TEMPERATURES

\underline{M. A. H. SMITH}, Science Directorate, NASA Langley Research Center, Hampton, VA 23681, U.S.A.; K. SUNG, L. R. BROWN, T. J. CRAWFORD, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr.,Pasadena, CA 91109, U.S.A.; A. W. MANTZ, Dept. of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320, U.S.A.; V. MALATHY DEVI, D. CHRIS BENNER, The College of William and Mary, Williamsburg, VA 23187, U.S.A..

High-resolution spectra of the ν_{4} fundamental band of ^{13}CH_{4} broadened by N_{2} at temperatures relevant to the atmosphere of Titan (80 K to 296 K) have been recorded using new temperature-controlled absorption cells installed in the sample compartment of a Bruker (IFS-125HR) Fourier Transform spectrometer (FTS) at the Jet Propulsion Laboratory (JPL). Details of the cells and spectrometer performance have been discussed in the previous talk.^{a} Early analysis of these spectra using multispectrum fitting has determined half widths, pressure-induced shifts, line mixing parameters and their temperature dependences for R-branch transitions from R(0) through R(6). In addition to the initial R(2) study mentioned in the previous talk, the analysis for the other J-manifolds examined in detail whether or not the N_{2}-broadened half width coefficients follow the simple power-law temperature-dependence over the entire temperature range from 80 K to 296 K. The results are compared with other published measurements of N_{2}-broadened methane parameters at low temperatures.^{b}

aA. W. Mantz et al., Closed-cycle He-cooled absorption cells designed for a Bruker IFS-125HR: First results between 79 K and 297 K, this session.

bResearch described in this paper was performed at Connecticut College, the College of William and Mary, NASA Langley Research Center and the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration.